At one large college, 1969 out of the 3500 students surveyed said they plan to take another science class. Find the 95% confidence interval for the proportion of students that said they plan to take another science class.

_?_% to _?_%
Answer #1

We know that 95% confidence interval for true proportion is given by:

\\(p\pm Z_{0.05/2}\sqrt{\frac{p(1-p)}{n}})\\ \\ (p\pm Z_{0.025}\sqrt{\frac{p(1-p)}{n}})\\ \\ (\frac{1969}{3500}\pm (1.96)\sqrt{\frac{\frac{1969}{3500}(1-\frac{1969}{3500})}{3500}})\\ \\ (0.5626\pm 0.0164)\\ \\ (0.5462,0.5790)